1. A solution containing 4.50 g of a non electrolyte (i=1) dissolved in 125 g of water





freezes at -0.372ºC. Calculate the molar mass of the solute.





Step 1: First compute the molality of the freezing point equation.





Step 2. Then, from the definition of molality, compute the number of moles





solute, n(solute), in the sample.





Step 3. Solve for the molar mass.









Please solve it like this







STEP1: SOLUTION (what formula you use and how you solve it)







STEP2:; SOLUTION (what formula you use and how you solve it)







STEP3: SOLUTION (what formula you use and how you solve it)

1
Expert's answer
2022-03-27T10:28:47-0400

"\Delta"Tf = i * Kf * m

"\Delta"Tf - change in freezing point

i - Van't Hoff factor = 1 (for non electrolyte)

Kf - freezing point constant of water = 1.86°C/m

m - molality of the solution

"\Delta"Tf = Tpure - Tsol = 0ºC - (-0.372ºC) = 0.372ºC

0.372ºC = 1.86°C/m * m

m = 0.372ºC / 1.86°C/m = 0.2 m

Molality = n(solute) / kg (solvent)

0.2 m = n(solute) / 0.125 kg

n(solute) = 0.125 kg * 0.2 m = 0.025 mol

n(solute) = m(solute) / Mr(solute)

Mr(solute) = m(solute) / n(solute) = 4.50 g / 0.025 mol = 180 g/mol